# Perturbative soft graviton theorems in de Sitter spacetime

### v2

@@ 6.2 Subleading and sub-subleading corrections
  • The gauge transformation at this order becomes very complicated in the momentum space. Thus, to check the gauge invariance of the subleading and sub-subleading corrections, one needs to check if eq. (54), from which they are derived, is gauge-invariant. Since eq. (54) is gauge invariant, the subleading and sub-subleading corrections are also gauge invariant.
@@
  • We have used a different set of scalar field modes (13) to determine the subleading and sub-subleading corrections to the flat space soft factor. We have observed that the subleading and sub-subleading corrections are exactly the same as eqs. (71) and (72) for this set of modes. Thus, the subleading and sub-subleading corrections are universal, and hence it may be related to asymptotic symmetries.
@@ 7 Discussions
In this paper, we obtained leading, subleading and sub-subleading order corrections induced by the de Sitter geometry in the large curvature length (or small cosmological constant) limit to the flat space soft graviton theorem, and discussed about the universal feature of these corrections. We first obtained mode solutions of the linearized field equation in the background de Sitter space. We then defined the inner product for these modes and verified that they are orthogonal. By defining the perturbative $\mathcal{S}$-matrix in a compact region of size $R$ inside the static patch and using the LSZ formula, we derived the leading, subleading, and sub-subleading corrections to the flat space soft factor. Further, consistency checks have been performed on these corrections. Performing the gauge transformations, we have verified that the leading, subleading, and sub-subleading corrections to the flat space soft factor are gauge-invariant. To check universality of these corrections, we used a different set of scalar field modes and observed that the leading order corrections depend on the choice of modes while the subleading and sub-subleading order corrections remain invariant. Thus, the leading corrections to the soft factor are non-universal, while the subleading and sub-subleading corrections are universal, which is consistent with the results obtained for soft photon theorems in Sayali_Diksha . The results are summarized in the following table-
@@
Flat space soft factor $\sum\limits_{i=1}^{n}\kappa\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}$ Universal Leading corrections $-im\sum\limits_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}\partial_{p_{i}}^{\beta}\bigg{)}$ Non-universal Subleading corrections $-m^{2}\sum\limits_{i=1}^{n}\kappa\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{3}}$ Universal Sub-subleading corrections $-m^{2}\sum\limits_{i=1}^{n}\kappa\bigg{(}\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{m^{2}(p_{i}\cdot\hat{k})^{2}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{3}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\partial_{p_{i}}^{\beta}\bigg{)}$ Universal
@@
Since the subleading and sub-subleading corrections are universal, it would be interesting to investigate whether perturbative corrections to the flat space asymptotic charges $(\mathcal{Q}_{l})$ can reproduce these universal terms. It is natural to expect such perturbative corrections should start at $\mathcal{O}(l^{-2})$. Recently, the soft graviton theorem in the de Sitter space was derived from near horizon supertranslation charges $(\mathcal{Q}_{H})$ Mao_Zhou , where the supertranslation charges were defined at the cosmological horizons. This raises the question of whether the near horizon charges $\mathcal{Q}_{H}$ reduce to the charges $\mathcal{Q}_{l}$ in the limit of large curvature length in the static patch. If this is the case, it may be possible to relate our results with those of Mao_Zhou . We leave it for future investigation.
@@
In the pursuit of understanding the gravitational radiation and memory effects in the presence of a tiny cosmological constant, the analysis presented here might be useful. However, to establish a relation between the soft factors deduced here as a perturbative expansion in cosmological constant, and the (asymptotic) symmetries of the static patch of the de Sitter spacetime, one needs to first analyze the symmetries of the patch of the full de Sitter spacetime and the corresponding charges. Since the static patch only preserves a 4 dimensional subset of the full 10 dimensional de Sitter symmetries, the scattering amplitudes should also obey the subset of symmetries only Ashtekar_Bonga ; Emil_2021 . It would be interesting to see if a relation between the asymptotic charges within this patch, and the soft factors can be established by performing a suitable expansion of the charges around the flat space. This would also be necessary to understand the memory effects in this framework Compere_Hoque . Finally, one needs to explore the connection between the asymptotic symmetry group of full de Sitter spacetime, the $\Lambda$-$\text{BMS}_{4}$ group Compere_Ruzziconi_2019 ; Compere_Ruzziconi_2020 and the symmetries of the soft amplitudes considered here.
@@ Acknowledgements
D.N.S. and P.C. want to thank Sayali Bhatkar and Diksha Jain for useful discussions during the initial stages of this work. D.N.S. would also like to thank Alok Laddha for useful discussions and his valuable suggestions. S. B. acknowledges support by DST-SERB/ANRF through MATRICS grant MTR/2022/000170.

### v3

@@ 6.4 Consistency checks
  • The leading correction to the flat space soft graviton theorem, Eq. (89), is invariant under the gauge transformation $\varepsilon_{\alpha\beta}\to\varepsilon_{\alpha\beta}-i\xi_{\alpha}k_{\beta}-i\xi_{\beta}k_{\alpha}$. Since the gauge transformation at $\mathcal{O}(l^{-2})$ becomes very complicated in the momentum space, to check the gauge invariance of the subleading and sub-subleading corrections to the flat space soft factor, we need to analyze whether Eq. (61), from which they are derived, is gauge-invariant. Since Eq. (61) is gauge invariant, the subleading and sub-subleading corrections to the flat space soft factor are also gauge invariant.
  • To check the universal feature of the leading correction term, we have used a different set of scalar field modes (13). The new set of modes has no terms of $\mathcal{O}(l^{-1})$, so there are no leading corrections to the flat space soft factor for this set of modes. Since the leading correction to the soft factor depends on the choice of modes, it is non-universal.
@@
  • We have used a different set of scalar field modes (13) to compute subleading and sub-subleading corrections to the flat space soft factor. The detailed calculation is given in the Appendix (A). We have observed that the subleading and sub-subleading corrections for this set of modes are exactly the same as Eqs. (79) and (80). Further, these correction terms only depend on the momenta of the hard particles, soft energy, and the coupling constant. Thus, we expect that the subleading and sub-subleading corrections to the flat space soft factor should be universal.
@@ 7 Discussions
In this paper, we obtained leading, subleading and sub-subleading order corrections induced by the de Sitter geometry in the large curvature length (or small cosmological constant) limit to the flat space soft graviton theorem, and discussed about the universal feature of these corrections. We first obtained mode solutions of the linearized field equation in the background de Sitter space. We then defined the inner product for these modes and verified that they are orthogonal. By defining the perturbative $\mathcal{S}$-matrix in a compact region inside the static patch, and using the LSZ formula, we derived the leading, subleading, and sub-subleading corrections to the flat space soft factor. Further, consistency checks have been performed on these corrections. Performing the gauge transformations, we have verified that the leading, subleading, and sub-subleading corrections to the flat space soft factor are gauge-invariant. To comment on the universality of these corrections, we used a different set of scalar field modes to recompute the leading, subleading, and sub-subleading corrections. We observed that the leading order corrections depend on the choice of modes while the subleading and sub-subleading order corrections remain the same. Thus, the leading corrections to the soft factor are non-universal, while we expect that the subleading and sub-subleading corrections are universal, which is consistent with the results obtained for soft photon theorems in Sayali_Diksha . The leading, subleading, and sub-subleading corrections to the flat space soft graviton theorems for the tree-level scattering process are summarized in the table below:
@@
Flat space soft factor $\sum\limits_{i=1}^{n}\dfrac{\kappa}{2}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{\omega p_{i}\cdot\hat{k}}-i\sum\limits_{i=1}^{n}\dfrac{\kappa}{2}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}\hat{k}_{\gamma}J^{\beta\gamma}}{p_{i}\cdot\hat{k}}$
Leading corrections $m\sum\limits_{i=1}^{n}\dfrac{\kappa}{2}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}\hat{k}_{\gamma}J^{\beta\gamma}}{(p_{i}\cdot\hat{k})^{2}}$
Subleading corrections $\dfrac{3m^{2}}{2}\sum\limits_{i=1}^{n}\dfrac{\kappa}{2}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{\omega(p_{i}\cdot\hat{k})^{3}}$
Sub-subleading corrections $\sum\limits_{i=1}^{n}\dfrac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{2}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot\hat{k})^{3}}\hat{k}\cdot\partial_{p_{i}}+\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot\hat{k})^{2}}p_{i}\cdot\partial_{p_{i}}+\dfrac{m^{2}(p_{i}^{\alpha}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot\hat{k})^{2}}\partial_{p_{i}}^{\beta}\bigg)$
@@
If the subleading and sub-subleading corrections turn out to be universal, it would be interesting to investigate whether perturbative corrections to the flat space asymptotic charges can reproduce these corrections from Ward identities. It is natural to expect such perturbative corrections should start at $\mathcal{O}(l^{-2})$. Recently, the soft graviton theorem in the de Sitter space was derived from Ward identities using the near-horizon supertranslation charges Mao_Zhou . Their analysis was done near to the horizon and it is not close to the small compact region where our setup is confined. For the validity of our results, one can not extend the finite boundaries of this region to the cosmological horizons. Also, taking the small cosmological constant limit of the results of Mao_Zhou pushes the cosmological horizons to the null infinities in the Minkowski space. Thus, their results can not be expanded naively in $\mathcal{O}(1/l)$ and directly reduce to the flat space soft graviton theorems. In addition to this, the authors in Mao_Zhou considered the massless scalars, and the massless limit of scalars is not well defined in our analysis especially for the scalar field modes. Thus, it is not immediately clear to us how our results and those of Mao_Zhou can be related.
Moreover, one can study the radiative component of the low frequency gravitational waveforms in the de Sitter background in the small cosmological constant limit to derive the perturbative corrections to the classical soft graviton theorems. One can also evaluate the perturbative corrections to the trajectories of hard particles, and then take the classical limit of our results by replacing all the operators with their classical counterparts. Evaluating perturbative corrections to the classical soft graviton theorem and interpreting various correction terms is an interesting future direction.
@@
In the pursuit of understanding gravitational radiation and memory effects in the presence of a tiny cosmological constant, the analysis presented here might be useful. However, to establish a relation between the soft factors deduced here as a perturbative expansion in cosmological constant, and the (asymptotic) symmetries of the static patch of the de Sitter spacetime, one needs to first analyze the symmetries of the patch of the full de Sitter spacetime and the corresponding charges. Since the static patch only preserves a 4 dimensional subset of the full 10 dimensional de Sitter symmetries, the scattering amplitudes should also obey the subset of symmetries only Ashtekar_Bonga ; Emil_2021 . It would be interesting to see if a relation between the asymptotic charges within this patch, and the soft factors can be established by performing a suitable expansion of the charges around the flat space. This would also be necessary to understand the memory effects in this framework Compere_Hoque . Finally, one needs to explore the connection between the asymptotic symmetry group of full de Sitter spacetime, the $\Lambda$-$\text{BMS}_{4}$ group Compere_Ruzziconi_2019 ; Compere_Ruzziconi_2020 and the symmetries of the soft amplitudes considered here.
@@ Acknowledgements
D.N.S. and P.C. would like to thank Sayali Bhatkar and Diksha Jain for useful discussions during the initial stages of this work. D.N.S. would also like to thank Alok Laddha for useful discussions and his valuable suggestions. S. B. acknowledges support by DST-SERB/ANRF through MATRICS grant MTR/2022/000170.
Appendix A Perturbative corrections for alternate set of scalar modes
In this section, we use the the alternate set of scalar field modes (13) to compute the perturbative corrections to the flat space soft factor. Even though these modes and the corresponding states are non-orthogonal, it is very interesting to study the soft limit of the $\mathcal{S}$-matrix for this set of modes. The $\mathcal{O}(l^{-2})$ corrections to the $\mathcal{S}$-matrix for this new set of modes is given as follows:
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\kappa\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\bigg(\dfrac{4m^{4}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{4}}+\dfrac{2m^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}$
$+\dfrac{2(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}+\dfrac{3ic(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2m(p^{2}+m^{2})}+\dfrac{5c(p_{i}\cdot z^{\prime})(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2m(p^{2}+m^{2})}-\dfrac{(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2m^{2}(p^{2}+m^{2})}$
$-\dfrac{ic(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2m(p^{2}+m^{2})}+\dfrac{i(p_{i}\cdot z^{\prime})^{3}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{6m^{2}(p^{2}+m^{2})}+\dfrac{z^{\prime 2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}-\dfrac{icmz^{\prime 2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}$
$+\dfrac{iz^{\prime 2}(p_{i}\cdot z^{\prime})(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}-\dfrac{2im^{2}(p\cdot z)(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}-\dfrac{i(p\cdot z)(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}+\dfrac{m^{2}z^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})^{2}}$
$-\dfrac{z^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}+\dfrac{3i(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}+\dfrac{cm(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{p^{2}+m^{2}}+\dfrac{(p\cdot z^{\prime})(p_{i}^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}-\dfrac{(p\cdot p_{i})(z^{\prime\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}\bigg)$
$\cdot e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle.$ (90)
To simplify the above expression, we first replace all $z^{\prime}$ by $i\partial_{p}$, and then perform the integration by parts in $p$. Next, we perform the $z^{\prime}$-integral which yields the delta function. We finally integrate this delta function and take the soft limit and obtain
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\frac{\kappa}{2}\int d^{4}z\bigg(\dfrac{3m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{2}}+\dfrac{3ic(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2m(p_{i}\cdot k)}$
$+\dfrac{im^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{2(p_{i}\cdot k)^{3}}-\dfrac{2i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}-\dfrac{cm(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}-\dfrac{3i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{2(p_{i}\cdot k)^{2}}$
$+\dfrac{5c(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{2m(p_{i}\cdot k)}+\dfrac{(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)(k\cdot z)}{2(p_{i}\cdot k)^{2}}-\dfrac{(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{2}}{2m^{2}(p_{i}\cdot k)}-\dfrac{ic(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{2}}{2m(p_{i}\cdot k)}$
$+\dfrac{i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)^{3}}{6m^{2}(p_{i}\cdot k)}-\dfrac{z^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{4(p_{i}\cdot k)}-\dfrac{icmz^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)}+\dfrac{iz^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{4(p_{i}\cdot k)}-\dfrac{im^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}$
$+\dfrac{4i(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{p_{i}\cdot k}+\dfrac{2cm(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{p_{i}\cdot k}+\dfrac{(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{2(p_{i}\cdot k)}+\dfrac{m^{2}(z^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{4(p_{i}\cdot k)}-\dfrac{z^{\alpha}z^{\beta}\varepsilon_{\alpha\beta}}{4}\bigg)$
$\bigg(1-ik\cdot z-\dfrac{1}{2}(k\cdot z)^{2}+...\bigg)e^{-ip_{i}\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle.$ (91)
By identifying the terms correspond to $A^{(0)}$ and $A^{(1)}$ in the above expression, and separating them from the rest of the terms, we can rewrite Eq. (A) as
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\frac{\kappa}{2}\int d^{4}z\bigg(\dfrac{3m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{2}}-\dfrac{im^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{3}}$
$-\dfrac{3i(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{2(p_{i}\cdot k)^{2}}-\dfrac{im^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}+\mathcal{O}(k^{-1})\bigg)e^{-ip_{i}\cdot z}$
$\cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle+A^{(1)}\Gamma_{n}^{(1)}+A^{(0)}\Gamma_{n}^{(2)},$ (92)
where $\mathcal{O}(k^{-1})$ terms contribute beyond the sub-subleading order and are suppressed. Now, we replace all $z$ by $i\partial_{p_{i}}$ in the above expression and put the $\mathcal{O}(l^{-2})$ corrections to the flat space soft factor in the momentum space as follows:
$A^{(2)}=\sum_{i=1}^{n}\frac{\kappa}{2}\bigg($ $\dfrac{3m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{2}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot k)^{3}}k\cdot\partial_{p_{i}}+\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}p_{i}\cdot\partial_{p_{i}}$
$+\dfrac{m^{2}(p_{i}^{\alpha}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}\partial_{p_{i}}^{\beta}\bigg)+\mathcal{O}(k^{-1}).$ (93)
We must note that it is independent of the arbitrary constant $c$ present in the scalar field modes in (13). The subleading and sub-subleading corrections to the flat space soft factor for this new set of modes (13) are
$S^{(2,-1)}=\sum_{i=1}^{n}\frac{\kappa}{2}\dfrac{3m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2\omega(p_{i}\cdot\hat{k})^{3}},$ (94)
$S^{(2,0)}=\sum_{i=1}^{n}\frac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{2}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot\hat{k})^{3}}\hat{k}\cdot\partial_{p_{i}}+\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot\hat{k})^{2}}p_{i}\cdot\partial_{p_{i}}+\dfrac{m^{2}(p_{i}^{\alpha}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot\hat{k})^{2}}\partial_{p_{i}}^{\beta}\bigg).$ (95)